### 10 Mar 2026 v1

@@ Abstract
Ternary relation; Betweenness relation; Fuzzy betweenness relation; KM-fuzzy metric space; Four-point transitivity; Five-point transitivity. AMS Subject Clasifications: 54A40; 03E72
@@ 1 Introduction
The current research on the concept of betweenness relations is conducted within the framework of an axiomatic structure. A betweenness relation is a special kind of ternary relation that satisfies reflexivity, symmetry, antisymmetry, and transitivity, which is widely associated with various mathematical structures, such as posets [23, 28, 39], lattices [6, 28], metrics [29, 30], median algebras [3, 10], normed spaces [5, 29, 31], convex structures [4, 32], road systems [2], and data aggregations [21]. So far, there are many definitions for the betweenness relation. The reason for this is that the definition of the transitivity of the ternary relation varies greatly. For example, there are definitions based on four points [11, 12] and those based on five points [22, 35]. In 2019, Pérez-Fernández and De Baets [21] gave a definition of betweenness relation which satisfies symmetry in the end points, closure and end-point transitivity. After that, Zhang et al. [39] represented a betweenness relation in sense of [21] as a family of order relations and analyze the corresponding family of induced Alexandrov topologies. These definitions and conclusions of the betweenness relations are all based on a specific type of four-point transitivity. Why do the researchers choose this particular kind of transitivity, and what is the origin and connection of these four-point and five-point transitivities? We will provide a detailed explanation in the section of Preliminaries.
@@
The metric space is a type of important and closest-to-Euclidean-space abstract space and it has extensive applications both in theoretical research and in real life. In a metric space, all triples that satisfy the metric triangle identity condition can generate a betweenness relation. In 1942, Menger [18] introduced the concept of a statistical metric space. Then Schweizer and Sklar [24, 25] developed Menger’s statistical metric and introduced a similar concept of a probabilistic metric space, which refers to the distance between two points corresponding to a probability distribution function defined on the set of non-negative real numbers. Besides, Menger also presented the concept of betweenness in [18] and pointed out its possible application fields. Shortly afterwards, Wald [33] extended Menger’s triangular inequality and established a theory of betweenness, which showed that the properties of Wald’s betweenness theory is consistent with the metric betweenness axioms. However, since then this topic has received very little attention.
@@
Influenced by the development of the fuzzy set theory [34], various definitions of fuzzy metrics were formally introduced by Kramosil and Michálek [15], Grabiec[9], Morsi [19], George and Veeramani [8]. In 1975, Kramosil and Michalek proved that in a certain sense, the KM-fuzzy metric is equivalent to the probability measure. To make the definition of the KM-fuzzy metric more complete, in 1988, Grabiec [21] added an another condition to the original definition of the KM-fuzzy metric and explained that the KM-fuzzy metric is not only a generalization of but also a semantic generalization of the classical metric. Subsequently, based on the monotonically decreasing fuzzy numbers, the Morsi’s fuzzy metric is introduced, which is equivalent to the complement of the KM-fuzzy metric. In 1994, George and Veeramani [8] modified some of the conditions in the definition of KM-fuzzy metrics, with the aim of making the induced topology Hausdorff. However, the topology induced by the GV-fuzzy metric is generated by the classical topology and cannot induce a fuzzy topology. In 2015, Mardones-Pérez and Prada Vicente [17] proposed another variant of the KM-fuzzy metric based on triangular norm operations. This new definition not only extended the original KM-fuzzy metric but also enabled it to induce multiple types of fuzzy structures. Morever, they discussed the characteristics of convergence, compactness, and completeness in KM-fuzzy pseudo-metric spaces.
@@
Recently, some scholars have once again paid attention to the research on the fuzzy betweenness relations. In 2020, Zhang, Pérez-Fernández, and De Baets [40] introduced two different definitions of fuzzy betweenness relations, which are *-betweenness relations and *-E betweenness relations. Soon afterwards, they investigated the issue of how to construct a fuzzy betweenness relation from a crsip metric with continuous Archimedean triangular norm in [41]. The description of fuzzy transitivity in these two definitions is based on the classic betweenness relation definition given by Pérez-Fernández, and De Baets [21] and it is also a fuzzy generalization of a certain specific type of the four-point transitive properties. In 2021, Zedam and De Baets [37] proposed the six types of transitivity properties of fuzzy ternary relations based on five points from the perspective of compositions. In 2022, Shi [26] explored the concept of gated sets within the framework of crisp betweenness relations induced by GV-fuzzy metric spaces. Also, Shi et al. [27] discussed the relationships between fuzzy betweenness relations and fuzzy interval operators, fuzzy partial orders and fuzzy Peano-Pasch spaces. In 2025, Jin and Yan provided a method to construct a *-betweenness relation and *-E betweenness relation by the GV-fuzzy metric and strong GV-fuzzy metric, respectively. However, through the above introduction, we can see that there are still many gaps in the current research on fuzzy betweenness relations by fuzzy metrics. By introducing the development status of the fuzziness metric, we can see that the KM-fuzzy metric is more in line with the semantic extension of fuzziness compared to the GV-fuzzy metric. This is also the starting point of our article. From the perspective of KM-fuzzy metric, we explore the the constructions and the characteristics of both betweenness relations and fuzzy betweenness relations in KM-fuzzy metric spaces.
@@
The main purpose of this paper is to introduce two different construction methods for fuzzy betweenness relations induced by a KM-fuzzy metric spaces. One approach is achieved by directly using the implication operator. The other method is realized through the corresponding nest of metrics generated by the KM-fuzzy metric. Furthermore, we demonstrate that these two types of fuzzy betweenness relations satisfy the eight kinds of four-point transitivity properties and the six kinds of five-point transitivity properties.

### 21 Mar 2026 v2

@@ Abstract
Ternary relation; Four-point transitivity; Five-point transitivity; Betweenness relation; Fuzzy betweenness relation; KM-fuzzy metric space. AMS Subject Clasifications: 54A40; 03E72
@@ 1 Introduction
The current research on the concept of betweenness relations is conducted within the framework of an axiomatic structure. A betweenness relation is a special kind of ternary relation that satisfies reflexivity, symmetry, antisymmetry, and transitivity, which is widely associated with various mathematical structures, such as posets [23, 28, 39], lattices [6, 28], metrics [29, 30], median algebras [3, 10], normed spaces [5, 29, 31], convex structures [4, 32], road systems [2], and data aggregations [21]. So far, there are many definitions for the betweenness relation. The reason for this is that the definition of the transitivity of the ternary relation varies greatly. For example, there are definitions based on four points [11, 12] and those based on five points [22, 35]. In 2019, Pérez-Fernández and De Baets [21] gave a definition of betweenness relation which satisfies symmetry in the end points, closure and end-point transitivity. After that, Zhang et al. [39] represented a betweenness relation in sense of [21] as a family of order relations and analyze the corresponding family of induced Alexandrov topologies. These definitions and conclusions of the betweenness relations are all based on a specific type of four-point transitivity.
@@
The metric space is a type of important and closest-to-Euclidean-space abstract space and it has extensive applications both in theoretical research and in real life. In a metric space, all triples that satisfy the metric triangle identity condition can generate a betweenness relation. In 1942, Menger [18] introduced the concept of a statistical metric space. Then Schweizer and Sklar [24, 25] developed Menger’s statistical metric and introduced a similar concept of a probabilistic metric space, which refers to the distance between two points corresponding to a probability distribution function defined on the set of non-negative real numbers. Besides, Menger also presented the concept of betweenness in [18] and pointed out its possible application fields. Shortly afterwards, Wald [33] extended Menger’s triangular inequality and established a theory of betweenness, which showed that the properties of Wald’s betweenness theory is consistent with the metric betweenness axioms. Since then, this topic has received very little attention.
@@
Influenced by the development of the fuzzy set theory, various definitions of fuzzy metrics were introduced by Kramosil and Michálek, Grabiec, Morsi, George and Veeramani, Mardones-Pérez and Prada Vicente. In 1975, Kramosil and Michalek [15] proved that in a certain sense, the KM-fuzzy metric is equivalent to the probability metric. To make the definition more complete, Grabiec [9] added an another condition to the original definition of the KM-fuzzy metric and explained the KM-fuzzy metric is not only a generalization but also a semantic generalization of the classical metric. Subsequently, based on the monotonically decreasing fuzzy numbers, the Morsi’s fuzzy metric [19] is introduced, which is equivalent to the complement of the KM-fuzzy metric. In 1994, George and Veeramani [8] modified some conditions, with the aim of making the induced topology Hausdorff. However, the topology induced by the GV-fuzzy metric is the classical topology and cannot induce a fuzzy topology. In 2015, Mardones-Pérez and Prada Vicente [17] proposed another variant of the KM-fuzzy metric through triangular norm operators. This new definition not only extended the original definition of KM-fuzzy metric but also enabled it to induce fuzzy structures.
@@
Recently, some scholars have once again paid attention to the research on the fuzzy betweenness relations. In 2020, Zhang, Pérez-Fernández, and De Baets [40] introduced two different definitions of fuzzy betweenness relations, which are *-betweenness relations and *-E betweenness relations. Soon afterwards, they investigated the issue of how to construct a fuzzy betweenness relation from a crsip metric with continuous Archimedean triangular norm in [41]. The description of fuzzy transitivity in these two definitions is based on the classic betweenness relation definition given by Pérez-Fernández, and De Baets [21] and it is also a fuzzy generalization of a certain specific type of the four-point transitive properties. In 2021, Zedam and De Baets [37] proposed the six types of transitivity properties of fuzzy ternary relations based on five points from the perspective of compositions. In 2022, Shi [26] explored the concept of gated sets within the framework of crisp betweenness relations induced by GV-fuzzy metric spaces. Also, Shi et al. [27] discussed the relationships between fuzzy betweenness relations and fuzzy interval operators, fuzzy partial orders and fuzzy Peano-Pasch spaces. In 2025, Jin and Yan provided a method to construct a *-betweenness relation and *-E betweenness relation by the GV-fuzzy metric and strong GV-fuzzy metric, respectively. However, through the above introduction, we can see that there are still many gaps in the current research on fuzzy betweenness relations by fuzzy metrics. By introducing the development status of the fuzziness metric, we can see that the KM-fuzzy metric is more in line with the semantic extension of fuzziness compared to the GV-fuzzy metric. This is also the starting point of our article. we will explore the the constructions and the characteristics of both betweenness relations and fuzzy betweenness relations in KM-fuzzy metric spaces.
@@
The main purpose of this paper is to introduce two different construction methods for fuzzy betweenness relations induced by a KM-fuzzy metric. One approach is achieved by directly using the implication operator. The other method is realized through the corresponding nest of metrics. Furthermore, we demonstrate that these two types of fuzzy betweenness relations satisfy the eight kinds of four-point transitivity properties and the six kinds of five-point transitivity properties.